Championship Boxing, Inc. is a small manufacturer of cardboard boxes of all sizes. You have reported for your first day of work, and the company is in an uproar. Yearly financial statements are being prepared, but a computer malfunction of the company’s new BOX-9000 computer has inadvertently erased parts of the company’s balance sheet, along with almost all related data except the company’s statement of cash flows. The IT department is working to retrieve earlier backups, but estimates that the reconstruction of the data will take about 24 hours.

Unfortunately, financial statements are to be presented at a stockholders’ meeting in one hour. The company uses the indirect method to prepare its statement of cash flows (rather than the direct method), so your new supervisor believes the missing data for the balance sheet can be prepared using the statement of cash flows. You are assigned this task, since you were top student in your business school class. Meanwhile, the supervisor will go to the stockholders’ meeting and give some introductory remarks.

In addition to the statement of cash flows, the following data survived the computer mishap:

A. The investments were sold for $280,000 cash.
B. Equipment was acquired for $152,000 cash.
C. Land was acquired for $326,000 cash.
D. There were no disposals of equipment during the year.
E. 12,500 shares of common stock were sold for cash during the year.
F

There was a $96,000 debit to Retained Earnings for cash dividends declared.

Your supervisor has provided you with the following statement of cash flows, prepared using the indirect method. Recall that the statement of cash flows consists of three sections: cash flows from operating activities, cash flows from investing activities, and cash flows from financing activities. Review the statement, and then proceed to the next panel.

Championship Boxing, Inc.

Statement of Cash Flows

For the Year Ended December 31, 20Y8

1

Cash flows from operating activities:

2

Net income

$186,540.00

3

Adjustments to reconcile net income to net cash flow from operating activities:

4

Depreciation

18,400.00

5

Gain on sale of investments

(50,000.00)

6

Changes in current operating assets and liabilities:

7

Increase in accounts receivable

(25,390.00)

8

Increase in inventories

(33,450.00)

9

Increase in accounts payable

41,130.00

10

Decrease in accrued expenses payable

(12,470.00)

11

Net cash flow from operating activities

$124,760.00

12

Cash flows from investing activities:

13

Cash received from sale of investments

$280,000.00

14

Cash paid for purchase of land

(326,000.00)

15

Cash paid for purchase of equipment

(152,000.00)

16

Net cash flow used for investing activities

(198,000.00)

17

Cash flows from financing activities:

18

Cash received from sale of common stock

$187,500.00

19

Cash paid for dividends

(91,200.00)

20

Net cash flow from financing activities

96,300.00

21

Change in cash

$23,060.00

22

Cash at the beginning of the year

585,920.00

23

Cash at the end of the year

$608,980.00

Do This Part:

Using the information on previous panels, complete the following comparative balance sheet.

Score: 35/76

Championship Boxing, Inc.

Comparative Balance Sheet

December 31, 20Y8 and 20Y7

1

Dec. 31, 20Y8

Dec. 31, 20Y7

2

Assets

3

Cash

$585,920.00

4

Accounts receivable (net)

230,950.00

5

Inventories

618,320.00

6

Investments

0.00

7

Land

0.00

8

Equipment

705,120.00

9

Accumulated depreciation-equipment

(166,400.00)

10

Total assets

11

12

Liabilities

13

Accounts payable (merchandise creditors)

$391,800.00

14

Accrued expenses payable (operating expenses)

41,150.00

15

Dividends payable

19,200.00

16

Total liabilities

$498,080.00

17

18

Stockholders’ Equity

19

Common stock, $4 par

100,000.00

20

Paid-in capital in excess of par

280,000.00

21

Retained earnings

1,290,840.00

22

Total stockholders’ equity

$1,858,340.00

23

Total liabilities and stockholders’ equity
Answer #1

The comparative balance sheet will be prepared as under:

Championship Boxing, Inc.
Comparative Balance Sheet
December 31, 20Y8 and 20Y7
1 Dec. 31, 20Y8 Dec. 31, 20Y7
2 Assets
3 Cash 6,08,980.00 5,85,920.00
4 Accounts receivable (net) 2,30,950.00 2,05,560.00
5 Inventories 6,51,770.00 6,18,320.00
6 Investments 0    2,30,000.00
7 Land 326000 0
8 Equipment 7,05,120.00 5,53,120.00
9 Accumulated depreciation-equipment -1,66,400.00 -1,48,000.00
10 Total assets 23,56,420.00 20,44,920.00
11
12 Liabilities
13 Accounts payable (merchandise creditors) 4,32,930.00 3,91,800.00
14 Accrued expenses payable (operating expenses) 41,150.00 53,620.00
15 Dividends payable 24,000.00 19,200.00
16 Total liabilities 4,98,080.00 4,64,620.00
17
18 Stockholders’ Equity
19 Common stock, $4 par 1,50,000.00 1,00,000.00
20 Paid-in capital in excess of par 4,17,500.00 2,80,000.00
21 Retained earnings 12,90,840.00 12,00,300.00
22 Total stockholders’ equity 18,58,340.00 15,80,300.00
23 Total liabilities and stockholders’ equity 23,56,420.00 20,44,920.00

Notes

Cash ending balance of $608,980 available from Cash Flow statement

AR beginning balance = Ending AR balance - increase in AR = 230950-25390 = 205560

Inventories ending balance = Beg. balance + increase in inventories = 618320 + 33450 = 651770

Investments balance = cash received - gain on sale of investments = 280000-50000 = 230000

Land = land purchased = 326000

Equipment beg. balance = ending balance - equipment purchased = 705120-152000 = 553120

Acc. Dep. beg. balance = ending bal. - depreciation expense = 166400-18400 = 148000

Total Assets will be the sum of all the above balances for the respective years.

Accounts Payable ending balance = beg. balance + increase = 391800 + 41130 = 432930

Accrued exp. beg. balance = ending balance + decrease = 41150+12470 = 53620

Total liabilities = sum of AP, accrued expenses and dividends payable

Retained Earnings beg. balance = Ending bal. - net income + dividends =